Session: Nov 2004 (session 3)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Xinyong Huang at Wed Nov 3 11:56:39 2004) Success :
We recorded the part data of the ea030 and the data of the ec019 together in the disk-pack SHAO-007. The scan 308-0634 was lost because of the power break off.

Medicina (Andrea Orlati at Wed Nov 3 12:50:38 2004) Minor failures : POIN
Communication link between station computer and antenna control unit fell down. This caused scans from 13 UT to 14:30 were lost.

Onsala (Michael Lindqvist at Wed Nov 3 12:51:11 2004) Success : HIGH
No known problems. High Tsys due to bad weather.

Urumqi (uao at Wed Nov 3 13:49:14 2004) Minor failures : PARI HIGH
no p-cal.

Urumqi (wang zhen at Wed Nov 3 13:51:21 2004) Minor failures : PARI HIGH
MKIV recorder,no p-cal.

Noto (P.Cassaro at Wed Nov 3 15:54:29 2004) Minor failures : HIGH
From 01:00 ut to the end of the experiment Tsys high due to very bad whether.

Effelsberg (A. Kraus at Thu Nov 4 10:30:28 2004) Success :

Jodrell_Bank (Alastair Gunn at Mon Nov 8 11:34:19 2004) Failure :
MkII was used. Mk2 could not observe due to ongoing engineering. No alternative 22GHz antenna available.

Cambridge (Alastair Gunn at Mon Nov 8 11:39:25 2004) Minor failures :
Missed scans from 14:25 to 14:35 UT due to Mk5 communication problem. No other known problems.

Metsahovi (Ari Mujunen at Tue Nov 9 14:11:52 2004) Minor failures : PCAL
No phase cal (no pcal in N04K2 either) as the box is being constructed, otherwise seemed nominal.